OpenStudy (anonymous):

whats R? A 3m south of east B 3m north C 3m 30 degrees north east D 3m 50 degrees south west

OpenStudy (anonymous):

what's the general question ? ^_^

OpenStudy (anonymous):

whats R? Magnitude of R and direction of R.

OpenStudy (anonymous):

how can we determine the magnitude of R and direction of R if there's no problem lol?

OpenStudy (anonymous):

i can show you a sample of a previous problem and how it's done.

OpenStudy (anonymous):

Nikolas Oman April 17 at 6:46pm best way to resolve each vector into perpendicular components. i = positive horizontal direction (to the right) j = positive vertical direction (up) A = -5 cos 65 i + 5 sin 65 j B = 5 cos 50 i - 5 sin 50 j C = -5 cos 45 i - 5 sin 45 j R = A + B + C = (-5 cos 65 + 5 cos 50 - 5 cos 45) i + (5 sin 65 - 5 sin 50 - 5 sin 45) j R = -2.435 i - 2.834 j Now use the Pythagorean Theorem to get the magnitude: Magnitude of R = √(2.435² + 2.834²) = 3.7 m Direction of R = arctan(-2.834/-2.435) = 229.3° ... so that's 49.3° south of west

OpenStudy (anonymous):

that's the question? :)

OpenStudy (anonymous):

no thats another problem. question is whats the magnitude and direction of R if vectors are: A 3m south of east B 3m north C 3m 30 degrees north east D 3m 50 degrees south west

OpenStudy (anonymous):

LOL >_< my bad, I thought those were the choices ^^"

OpenStudy (anonymous):

lol no do you know how to do it?

OpenStudy (anonymous):

I'll give it a try? :)

OpenStudy (anonymous):

tnx :)

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

remind me, In the previous problem, A = -5 cos 65 i + 5 cos65 j < -- angles were given? or is that the general form

OpenStudy (anonymous):

angles were given

OpenStudy (anonymous):

how about for this problem?

OpenStudy (anonymous):

A 3m south of east B 3m north C 3m 30 degrees north east D 3m 50 degrees south west

OpenStudy (anonymous):

this is all that's given

OpenStudy (anonymous):

hmm, that means that you have to figure out the angles from the given

OpenStudy (anonymous):

B is easy, 3 m north , angle = 90 degrees

OpenStudy (anonymous):

hmm yea and A? 45 degrees?

OpenStudy (anonymous):

since they haven't given you the angles, yes you can assume that A = 45 degrees

OpenStudy (anonymous):

and now? how much is R?

OpenStudy (anonymous):

all you've got to do is use the same way you've used in the previous question :) I think your prob was trying to figure out the angles? right?

OpenStudy (anonymous):

well it's not my homework :) it's my friends homework and I have no idea how to do it :)

OpenStudy (anonymous):

lol >_<

OpenStudy (anonymous):

Alright, let's try to solve it together ^_^

OpenStudy (anonymous):

she solved it and got 2.79m 1.65 degrees north of east and she just wants to check if it's ok

OpenStudy (anonymous):

so: i = positive horizontal direction (to the right) j = positive vertical direction (up) then we can say the following : A = 3 cos 45 - 3 sin 45 B = 3 sin 90 C = 3 cos 30 + 3 sin 30 D = - 3 cos 50 - 3 sin 50

OpenStudy (anonymous):

yep i agree so far

OpenStudy (anonymous):

now: R = A + B + C = (3 cos 45 - 3 sin 45 ) + (3 sin 90 ) + (3 cos 30 + 3 sin 30) + (-3 cos 50 - 3 sin 50) = 0 + 3 + pi - 4.23 = 1.912

OpenStudy (anonymous):

R= 3 cos 45 + 3 cos 30 - 3 cos 50)i + (3 sin 45 + 3 sin 90 + 3 sin 30 - 3 sin 50)j

OpenStudy (anonymous):

lol, wait, I've messed it up

OpenStudy (anonymous):

yeah, hold on :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

R = 1.76 i - 3.32 j I might have some calculation mistakes, it happens ^_^

OpenStudy (anonymous):

so this is it?

OpenStudy (anonymous):

\[R = \sqrt((1.76)^2 + (3.32)^2)\] = 3.74 m

OpenStudy (anonymous):

alright thank you so very much!

OpenStudy (anonymous):

arctan = -62.07

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

thank you thank you so her answer was wrong?

OpenStudy (anonymous):

I'm not sure what was her answer lol , I just followed the steps and given :)

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

np ^_^, but I'm not sure though ^^"

OpenStudy (anonymous):

is there an online calculator where i could check this?

OpenStudy (anonymous):

there's your hand-calculator? :)

OpenStudy (anonymous):

i dont have it :/

OpenStudy (anonymous):

can you calculate this for me please? R = (3 cos 45 + 3 cos 30 - 3 cos 50)i + (-3 sin 45 + 3 sin 90 + 3 sin 30 - 3 sin 50)

OpenStudy (anonymous):

I did lol, that's what I got, answer above

OpenStudy (anonymous):

you can check with your calculator if you want :)

OpenStudy (anonymous):

oh ok thank you again :)

OpenStudy (anonymous):

lol, np ^_^